OpenStudy (loser66):

compute \(2^{2r}\) (mod 3), for all positive integers r. Please, check

OpenStudy (loser66):

\(2^{2r}=4^r\) and \(4\equiv 1(mod3)\) so that \(2^{2r}\equiv 1(mod3)\) Am I right?

OpenStudy (amistre64):

seems fair to me.

OpenStudy (loser66):

Please, help on this: Let t be an integers, and let n be a positive integer. Prove that: \[Z_n= \{[0],[1],....,[n-1]\}=\{[t],[t+1],....,[t+(n-1)]\}\] How to start?

OpenStudy (amistre64):

do the square brackets mean anything special?

OpenStudy (loser66):

That is class in modulo n

OpenStudy (loser66):

just terminology, like in Z_5, 4+1= 0, but we use [4]+[1]=0

OpenStudy (amistre64):

Z mod n is defined to be the set of classes ..... got it

OpenStudy (amistre64):

im thinking the proof has something to do with the cyclic nature of the group

OpenStudy (amistre64):

by convention, we organize it 0 to n-1 but what we name the elements is of little consequence to the properties of the group

OpenStudy (amistre64):

since t is some element within a class of Z mod n, and the name of the set can be referencned by any element of the set ... something along those lines

OpenStudy (loser66):

We don't have it, we need to prove it. From the beginning, t is just an integer in Z

OpenStudy (amistre64):

just an idea, not to sure how to put it into 'proper' stuff tho :) but for example say t = 5 mod 8 the set Z8 = {5,6,7,8,9,10,11,12} = {5,6,7,0,1,2,3,4} = {0,1,2,3,4,5,6,7}

OpenStudy (loser66):

We have elements in Z_n is the middle term of the expression. I mean \(Z_n= \{[0],[1],....,[n-1]\}\) and we need to prove that if n|t, then the remainder will be \(\{[t],[t+1],....,[t+(n-1)]\}\)

OpenStudy (loser66):

Can we prove the group of the middle one isomophic to the group of the right one?

OpenStudy (amistre64):

thats along the lines of what im thinking yes. simply by showing the t is congruent to some class, [x] in Z mod n. and by definition t+1 is congruent to the class [x+1] etc

OpenStudy (amistre64):

in my example: 5 = 13 mod 8 just as well as 13 = 5 mod 8 because of symetry might be the proper name of for it

OpenStudy (loser66):

Continue, please. We didn't reach somewhere yet. :)

OpenStudy (amistre64):

since Z mod n is cyclic, and t = x mod n, then for some k, t+k = x+k = 0 mod n, but without my texts handy a proper write up wouldnt be possible for me

OpenStudy (amistre64):

but my proof attempt would be along those thoughts

OpenStudy (anonymous):

I'm not sure how much about Z (mod n) is standard knowledge, so some of the things I post might seem trivial. Claim 1: Let $$0\le i \le n-1.$$ If $$t+i=qn+r$$ for some integers q and r with: $$0\le r \le n-1,$$ then: $$[t+i]=[r].$$ Proof: If t+i=qn+r, then (t+i)-r=qn, a multiple of n. This is exactly the definition of what it means to be congruent mod n.

OpenStudy (anonymous):

Claim 2: \[\left\{ [t],[t+1], \ldots , [t+(n-1)] \right\}\subset \left\{ [0],[1], \ldots , [n-1] \right\}\] Proof: It turns out any integer b can be written as b = qn+r with $$0\le r\le n-1.$$This is the division algorithm. So for any i, [t+i]=[r] for some r between 0 and n-1. Now all thats left to do is show that each [t+i] is distinct. That is...

OpenStudy (anonymous):

Claim 3: Let $$0\le i,j \le n-1, i\ne j.$$ Then $$[t+i]\ne [t+j].$$ Proof: Either a proof by contradiction or the contrapositive will work here. I'll work by contradiction. If [t+i]=[t+j], then [t] + [i] = [t] + [j], whence [i] = [j]. This means i-j = qn for some q in Z. However, since both i and j are integers between 0 and n-1, their difference is smaller than n, i.e: $$|i-j|<n.$$ Therefore, if i-j=qn, q must be 0, and we see that i-j=0 which implies i=j, a contradiction.

OpenStudy (anonymous):

Now that we know all the [t+i] are distinct, its clear that:\[\left\{ [t],[t+1],\ldots , [t+(n-1)] \right\}=\left\{ [0],[1],\ldots , [n-1] \right\}\]since one is contained in the other, and they both have the same number of elements.

OpenStudy (freckles):

How weird would it be to do induction. So we can prove for all natural t we have \[\left\{ [t],[t+1],\ldots , [t+(n-1)] \right\}=\left\{ [0],[1],\ldots , [n-1] \right\} \] --- Clearly it works for t=0 And for t=1, we have [0] in that one set but not in the other or at least we have him in disguise in the other group as [n]. So for assume for some integer t=k we have \[\left\{ [k],[k+1],\ldots , [k+(n-1)] \right\}=\left\{ [0],[1],\ldots , [n-1] \right\} \] Now we want to show it is true for t=k+1 \[\left\{ [k+1],[k+2],\ldots , [k+n)] \right\}=\left\{ [0],[1],\ldots , [n-1] \right\} \] So only element that wasn't in set Pk that in P(k+1) is [k+n] and the only element that is in P(k+1) that is not in Pk is [k]. But [k] is equivalent to [n+k]. So that meants \[\left\{ [k],[k+1],\ldots , [k+(n-1)] \right\}=\left\{ [k+1],[k+2],\ldots , [k+n)] \right\}=\left\{ [0],[1],\ldots , [n-1] \right\} \]